# Distributed Harmonization: Federated Clustered Batch Effect Adjustment and Generalization

### v1

@@ 5.1. ADNI Data
We use neuroimaging data from the second phase of the North American Alzheimer’s Disease Neuroimaging Initiative (ADNI) to evaluate our proposed methods. The ADNI data we used has MRI imaging of 563 scans/subjects collected from 18 participating sites. We extracted regional measures from DTI data, following the procedure in (Nir et al., 2013), leading to 228 features from each scan. We also construct a set of downstream prediction tasks, including the prediction of a set of ADNI-defined indicators derived from the neuropsychological battery to characterize memory, executive function, and language. Specifically: 1) MEM: The ADNI-Mem composite score for memory, which is based on the Rey Auditory Verbal Learning task, word list learning and recognition tasks from ADAS-Cog, recall from Logical Memory I of the Wechsler Memory Test–Revised, and the 3-word recall item from the MMSE (Crane et al., 2012). 2) EXF: ADNI-EF composite score for executive function, including Category Fluency (i.e., animals and vegetables), Trail-Making Test parts A and B, Digit Span Backwards, Wechsler Adult Intelligence Scale–Revised Digit–Symbol Substitution, and 5 Clock Drawing items (Gibbons et al., 2012). 3) LAN: ADNI-Lan indicator, which is a composite measure of language (Gibbons et al., 2012). We also include changes in these scores from baselines (Hohman et al., 2017), denoted by MEM SLOPES, EXF SLOPES, LAN SLOPES, respectively. Later, we use these six target variables to evaluate regression performance in downstream tasks. The characteristic distribution of the ADNI dataset is illustrated in Table 3.
@@ 5.2. Site and Cluster Effects in Brain Imaging
For site effect visualization, we only show distributions of cluster 1 and cluster 4 for demonstration. And we color the same site-index-based LDA visualization  [^1] LDA visualization results will differ depending on the choice of target variable. using different coloring strategies, for a better understanding of relations between site and cluster effect in Cluster ComBat: the left column figures (Figure 6(a), 6(c)) are colored by site index, while the right column figures (Figure 6(b), 6(d)) are colored by cluster index. By comparing Figure 6(a) and 6(b), we can know that both site effect and cluster effect are evident in unharmonized data, as distinct separation is observed between sites and clusters. By comparing Figure 6(b) and 6(d), we verify that our Cluster ComBat does remove cluster effect, as cluster 1 and cluster 4 overlap with each other after harmonization. Then, by coloring samples in the same cluster differently based on site index, as shown in Figure 6(c), we find that cluster 1 consists of site 3, 6, 9 and 12. Though site 6 and 12 overlap with each other, site 3, 6 and 9 are clearly separated from each other. Similar in cluster 4, site 1 and 8 show obvious disparity with each other. To conclude, our Cluster ComBat removes differences over clusters while preserving possible site differences within the cluster, which is also verified in higher site classification accuracy than ComBat in Table 4.
@@ 5.3. Downstream Regression Performance
For real data, we do not have the ground truth of harmonization, so our focus is on evaluating the performance of harmonization algorithms through downstream tasks. In these tasks, we use the 228 features of DTI brain imaging to predict the MEM, MEM SLOPES, EXF, EXF SLOPES, LAN, LAN SLOPES variables. We build a simple Linear Regression model using the Scikit-Learn library (Pedregosa et al., 2011) to train the regression task on the target target variables. For ComBat and Distributed ComBat, we retrain parameters as described in Section 4.2. We split the 18 sites into 12 training sites and 6 testing sites, then run experiments 100 times with different combinations of train and test sites. To evaluate performance, we compute the Mean Absolute Error (MAE) of the linear regression’s outputs on the testing site’s data and target testing labels. In a centralized setting, we also compared our method with Generalized Linear Squares Approach (Wang et al., 2018), an algorithm designed to eliminate confounding effects. This approach assumes that a variable may be linearly dependent on the confounding variables, and these effects can be removed by solving a linear regression optimization problem. Results in Table 5 show that our proposed method performs better than ComBat and Generalized Linear Squares Approach in a centralized setting and Distributed ComBat in a decentralized setting for most downstream tasks.
@@ 5.4. Additional Empirical Studies
Important Feature Before and After Harmonization. For regression tasks, we compute $p$-values for linear regression across 228 features in DTI imaging. The final $p$-values are obtained by averaging over 100 different random seeds. A feature is important if its $p$-value is less than 0.05. Table  9 displays the number of important features for the linear regression across 3 target variables MEM, EXF, and LAN. The table indicates that, by using Cluster ComBat, we achieve comparable performance with fewer significant features. This suggests that without harmonization and ComBat, the model may have included too many false positive features.
@@ 6. Discussion and Conclusion
We conducted extensive validation on both synthetic data real brain imaging data from ADNI in both centralized and decentralized settings. We showed that our methods do remove cluster-wise effects from brain imaging data via both qualitative and quantitative studies. Then, our methods exhibit superior performance on downstream regression tasks compared to baseline harmonization methods in both centralized and decentralized settings, which further validates the efficacy of our harmonized data. We also showed that our methods can use much fewer significant features to achieve similar regression performance compared with other baselines, suggesting potential avenues for further research on selected features.
Regarding deploying our proposed method, we consider 3 implementation consideration aspects: 1) ML-framework agnostic:Our algorithm doesn’t involve any specific ML frameworks in the local computation part, so it is easy to implement in many systems regardless of the local ML framework. Flower (Beutel et al., 2022) can be a candidate choice. For downstream tasks after harmonization, like regression or other ML models, the choice of local ML framework can be flexible depending on local preference.   2) Security communication:Designed for medical records, the deployment system needs to have communication security to prevent privacy leakage. One possible choice is to encrypt the communication between the clients and the server, for example, Secure Socket Layer (SSL) (Hidayat et al., 2023) or Transport Layer Security (TLS) (Jalali and Chen, 2024).   3) Scalable and light-weight:Since our algorithm’s main benefit lies in new clients joining the federated system, the system deployment should be scalable. To be more specific, when new clients join in, there should be minimum system configuration modification on the server as well as for old clients. Also, the implementation of our algorithm needs to be light-weight, and the FL system with our algorithm should require limited system consumption. And the design of FedLab (Zeng et al., 2023) can be a reference to meet these requirements. As a future work, we will deploy our proposed Cluster ComBat harmonization in the ENIGMA Consortium toolbox to further validate existing studies.
@@ 7. Acknowledgement
This material is based in part upon work supported by the National Science Foundation under Grant IIS-2212174, IIS-1749940, Office of Naval Research N00014-20-1-2382, and National Institute on Aging (NIA) RF1AG072449, R01AG051628, R01AG056102.

### v3

@@ 5.1. ADNI Data
We use neuroimaging data from the second phase of the North American Alzheimer’s Disease Neuroimaging Initiative (ADNI) to evaluate our proposed methods. The ADNI data we used has MRI imaging of 563 scans/subjects collected from 18 participating sites. We extracted regional measures from DTI data, following the procedure in (Nir et al., 2013), leading to 228 features from each scan. We also construct a set of downstream prediction tasks, including the prediction of a set of ADNI-defined indicators derived from the neuropsychological battery to characterize memory, executive function, and language. Specifically: 1) MEM: The ADNI-Mem composite score for memory, which is based on the Rey Auditory Verbal Learning task, word list learning and recognition tasks from ADAS-Cog, recall from Logical Memory I of the Wechsler Memory Test–Revised, and the 3-word recall item from the MMSE (Crane et al., 2012). 2) EXF: ADNI-EF composite score for executive function, including Category Fluency (i.e., animals and vegetables), Trail-Making Test parts A and B, Digit Span Backwards, Wechsler Adult Intelligence Scale–Revised Digit–Symbol Substitution, and 5 Clock Drawing items (Gibbons et al., 2012). 3) LAN: ADNI-Lan indicator, which is a composite measure of language (Gibbons et al., 2012). We also include changes in these scores from baselines (Hohman et al., 2017), denoted by MEM SLOPES, EXF SLOPES, and LAN SLOPES, respectively. Later, we use these six target variables to evaluate regression performance in downstream tasks. The characteristic distribution of the ADNI dataset is illustrated in Table 3.
@@ 5.2. Site and Cluster Effects in Brain Imaging
For site effect visualization, we only show distributions of cluster 1 and cluster 4 for demonstration. And we color the same site-index-based LDA visualization  [^1] LDA visualization results will differ depending on the choice of target variable. using different coloring strategies, for a better understanding of relations between site and cluster effect in Cluster ComBat: the left column figures (Figure 6(a), 6(c)) are colored by site index, while the right column figures (Figure 6(b), 6(d)) are colored by cluster index. By comparing Figure 6(a) and 6(b), we can know that both site effect and cluster effect are evident in unharmonized data, as distinct separation is observed between sites and clusters. By comparing Figure 6(b) and 6(d), we verify that our Cluster ComBat does remove cluster effect, as cluster 1 and cluster 4 overlap with each other after harmonization. Then, by coloring samples in the same cluster differently based on site index, as shown in Figure 6(c), we find that cluster 1 consists of site 3, 6, 9 and 12. Though site 6 and 12 overlap with each other, site 3, 6 and 9 are clearly separated from each other. Similar to cluster 4, site 1 and 8 show obvious disparity with each other. To conclude, our Cluster ComBat removes differences over clusters while preserving possible site differences within the cluster, which is also verified in higher site classification accuracy than ComBat in Table 4.
@@ 5.3. Downstream Regression Performance
For real data, we do not have the ground truth of harmonization, so our focus is on evaluating the performance of harmonization algorithms through downstream tasks. In these tasks, we use the 228 features of DTI brain imaging to predict the MEM, MEM SLOPES, EXF, EXF SLOPES, LAN, and LAN SLOPES variables. We build a simple Linear Regression model using the Scikit-Learn library (Pedregosa et al., 2011) to train the regression task on the target target variables. For ComBat and Distributed ComBat, we retrain parameters as described in Section 4.2. We split the 18 sites into 12 training sites and 6 testing sites, then run experiments 100 times with different combinations of train and test sites. To evaluate performance, we compute the Mean Absolute Error (MAE) of the linear regression’s outputs on the testing site’s data and target testing labels. In a centralized setting, we also compared our method with the Generalized Linear Squares Approach (Wang et al., 2018), an algorithm designed to eliminate confounding effects. This approach assumes that a variable may be linearly dependent on the confounding variables, and these effects can be removed by solving a linear regression optimization problem. Results in Table 5 show that our proposed method performs better than ComBat and Generalized Linear Squares Approach in a centralized setting and Distributed ComBat in a decentralized setting for most downstream tasks.
@@ 5.4. Additional Empirical Studies
Important Feature Before and After Harmonization. For regression tasks, we compute $p$-values for linear regression across 228 features in DTI imaging. The final $p$-values are obtained by averaging over 100 different random seeds. A feature is important if its $p$-value is less than 0.05. Table  9 displays the number of important features for the linear regression across 3 target variables MEM, EXF, and LAN. The table indicates that by using Cluster ComBat, we achieve comparable performance with fewer significant features. This suggests that without harmonization and ComBat, the model may have included too many false positive features.
@@ 6. Discussion and Conclusion
We conducted extensive validation on both synthetic data real brain imaging data from ADNI in both centralized and decentralized settings. We demonstrated through both qualitative and quantitative studies that our methods effectively remove cluster-wise effects from brain imaging data. Then, our methods exhibit superior performance on downstream regression tasks compared to baseline harmonization methods in both centralized and decentralized settings, which further validates the efficacy of our harmonized data. We also showed that our methods can use much fewer significant features to achieve similar regression performance compared with other baselines, suggesting potential avenues for further research on selected features.
Regarding deploying our proposed method, we consider 3 implementation consideration aspects: 1) ML-framework agnostic:Our algorithm doesn’t involve any specific ML frameworks in the local computation part, so it is easy to implement in many systems regardless of the local ML framework. Flower (Beutel et al., 2022) can be a candidate choice. For downstream tasks after harmonization, like regression or other ML models, the choice of local ML framework can be flexible depending on local preference.   2) Security communication:Designed for medical records, the deployment system needs to have communication security to prevent privacy leakage. One possible choice is to encrypt the communication between the clients and the server, for example, Secure Socket Layer (SSL) (Hidayat et al., 2023) or Transport Layer Security (TLS) (Jalali and Chen, 2024).   3) Scalable and light-weight:Since our algorithm’s main benefit lies in new clients joining the federated system, the system deployment should be scalable. To be more specific, when new clients join in, there should be minimum system configuration modification on the server as well as for old clients. Also, the implementation of our algorithm needs to be lightweight, and the FL system with our algorithm should require limited system consumption. And the design of FedLab (Zeng et al., 2023) can be a reference to meet these requirements. As a future work, we will deploy our proposed Cluster ComBat harmonization in the ENIGMA Consortium toolbox to further validate existing studies.
@@ 7. Acknowledgement
This material is based in part upon work supported by the National Science Foundation under Grant IIS-2212174, IIS-1749940, IIS 2319450, IIS 2045848, Office of Naval Research N00014-24-1-2168, and National Institute on Aging (NIA) RF1AG072449, U01AG068057, National Institute of Mental Health RF1MH125928.